# SPORTU: A Comprehensive Sports Understanding Benchmark for Multimodal Large Language Models

### v2

@@ B.2 Results Across Different Levels of Difficulty and Sports for Each Model
Model(%) American Football Badminton Baseball Basketball Ice Hockey Soccer Volleyball Overall Easy Claude-3.0-Haiku 76.77 41.45 93.10 73.33 75.00 62.77 85.25 68.41 Claude-3.5-Sonnet 92.69 68.16 98.89 96.14 90.87 88.50 100.00 88.65 Gemini 1.5 Pro 86.14 79.91 96.45 79.23 76.98 86.87 100.00 85.85 Gemini 1.5 Flash 83.48 77.56 97.12 84.56 81.75 86.59 100.00 85.99 GPT-4omini 73.06 39.96 93.33 67.37 69.84 60.24 100.00 66.09 GPT-4o 87.90 68.16 96.44 91.58 82.54 84.06 100.00 84.89 GPT-4V 89.50 62.61 98.00 91.93 78.97 82.61 100.00 83.71 ChatUniVi 67.65 24.15 93.79 55.09 69.05 54.89 68.85 59.22 LLaVA-NeXT 96.12 82.26 99.56 98.59 96.03 94.66 98.36 94.24 mPLUG-Owl3 86.07 84.83 97.34 86.27 82.54 85.33 100.00 87.28 Tarsier 88.86 82.48 98.23 94.04 83.73 89.31 100.00 89.71 Video-ChatGPT 67.88 40.17 92.87 54.74 66.67 59.42 39.34 62.36 VideoChat2 90.68 72.22 96.90 94.01 89.29 88.95 88.52 88.30 Medium Claude-3.0-Haiku 44.08 49.36 43.78 50.15 53.83 45.57 45.02 46.43 Claude-3.5-Sonnet 66.00 74.13 85.80 73.04 52.51 55.04 68.25 65.10 Gemini 1.5 Pro 62.45 73.77 79.63 55.12 50.92 47.41 61.14 58.25 Gemini 1.5 Flash 52.19 56.47 45.43 57.38 56.99 54.35 52.61 53.38 GPT-4omini 54.61 84.44 75.00 54.97 46.70 49.65 40.28 55.69 GPT-4o 63.73 85.33 81.92 66.57 51.72 53.86 55.92 62.31 GPT-4V 53.25 69.22 76.64 66.42 59.63 53.65 47.39 59.79 ChatUniVi 31.69 35.34 25.76 37.95 49.87 41.61 29.86 36.95 LLaVA-NeXT 70.95 74.13 90.05 75.30 61.74 54.06 71.56 67.02 mPLUG-Owl3 52.94 76.50 71.66 54.37 45.38 48.51 52.13 55.58 Tarsier 56.99 58.83 68.85 68.07 62.00 52.14 57.82 58.43 Video-ChatGPT 45.95 40.25 35.61 36.30 41.42 39.12 32.23 39.40 VideoChat2 55.86 59.02 63.82 71.54 67.81 54.47 43.13 58.73 Hard Claude-3.0-Haiku 22.35 6.17 16.03 31.50 25.42 26.82 36.36 20.12 Claude-3.5-Sonnet 41.57 79.01 52.63 36.22 50.00 52.94 45.45 52.57 Gemini 1.5 Pro 41.34 78.40 64.08 24.60 43.22 40.24 45.45 53.53 Gemini 1.5 Flash 61.11 12.96 34.49 43.31 46.61 44.00 36.36 38.73 GPT-4omini 48.31 84.44 53.71 45.67 38.98 38.82 40.28 47.54 GPT-4o 39.89 78.40 67.98 37.80 56.78 44.94 36.36 57.84 GPT-4V 50.00 65.43 69.28 69.89 64.22 59.88 56.80 45.48 ChatUniVi 36.11 12.96 40.69 22.83 22.03 27.06 27.27 32.21 LLaVA-NeXT 37.64 9.88 32.46 45.24 51.69 33.88 36.36 33.44 mPLUG-Owl3 30.34 11.11 19.93 27.56 29.66 37.18 36.36 25.40 Tarsier 42.22 8.64 18.38 29.92 40.68 27.29 36.36 24.25 Video-ChatGPT 17.98 40.12 37.80 23.62 22.88 31.76 13.64 32.55 VideoChat2 36.31 6.79 20.17 37.01 22.88 41.88 31.82 26.94 All Levels Claude-3.0-Haiku 50.73 40.29 43.28 54.09 56.48 48.34 52.72 48.07 Claude-3.5-Sonnet 70.68 72.43 75.57 74.81 65.02 64.10 73.13 69.52 Gemini 1.5 Pro 66.67 76.84 77.09 57.91 58.48 57.59 68.03 65.13 Gemini 1.5 Flash 61.96 58.86 52.03 62.92 63.68 62.19 61.22 59.97 GPT-4omini 59.04 52.75 70.53 57.16 53.27 51.43 53.06 57.24 GPT-4o 67.77 74.62 79.52 69.80 62.88 61.29 63.61 68.00 GPT-4V 62.99 66.07 69.28 69.89 64.22 59.88 56.80 64.01 ChatUniVi 42.27 27.82 45.92 40.71 51.94 43.74 37.75 42.35 LLaVA-NeXT 74.20 68.53 69.53 77.93 71.70 63.17 74.49 68.89 mPLUG-Owl3 59.62 70.82 56.84 59.63 55.41 57.52 60.88 59.26 Tarsier 64.20 61.32 55.30 70.45 65.95 59.80 64.97 61.32 Video-ChatGPT 48.92 40.20 48.50 39.68 46.99 43.97 32.31 44.63 VideoChat2 63.37 57.08 53.71 73.40 67.96 62.69 51.70 61.55
B.3 Criteira of G-Eval
This score assesses the overall accuracy, conciseness, and relevance of the model-generated explanation to the ground truth explanation. The explanation should focus on identifying and clearly explaining the key issue or relevant action.
Scoring Breakdown:
  • Score 1 (very poor):
    The explanation is incorrect or does not address the key issue at all. It might be filled with irrelevant details that distract from the main point.
  • Score 2 (poor):
    The explanation contains some correct elements but fails to clearly identify the key issue. There may be excessive irrelevant details that obscure the main point.
  • Score 3 (adequate):
    The explanation identifies the key issue but includes too much unnecessary information or lacks clarity. The key point is present but could be more concise.
  • Score 4 (good):
    The explanation clearly identifies the key issue with some minor unnecessary details. The core explanation is correct and relevant but may include a few extra, non-essential details.
  • Score 5 (excellent):
    The explanation is concise, accurate, and directly addresses the key issue without any unnecessary information. It clearly and effectively answers the question.
Evaluation Steps:
  1. 1.
    Read the ground truth explanation.
  2. 2.
    Identify the specific context in the ground truth explanation that represents the key issue (e.g., a foul, a mistake).
  3. 3.
    Read the model output explanation.
  4. 4.
    Compare the model’s explanation to the ground truth explanation, focusing on how directly and concisely the model output identifies the key issue.
  5. 5.
    Assign a score for explanation of overall relevance on a scale of 1 to 5, based on the criteria.
Model Generated Explanation: {{Model Generated Content}}
Ground Truth Explanation: {{Ground Truth Content}}
Evaluation Form (scores ONLY): Overall:
B.4 Correlation Matrix of Evaluation Metrics
@@
We applied Pearson correlation to assess the relationship between the automatic evaluation metrics and human-annotated scores. As shown in Figure 4, G-eval has the highest correlation with the human scores among all the automatic metrics. However, the correlation is still relatively weak, indicating the need for the development of a new evaluation metric in the future to improve the assessment process
Figure 4: Correlation matrix of different evaluation metrics in SPORTU-video
B.5 Prompt Templates
Detailed templates for zero- and five-shot standard prompts, as well as zero- and five-shot CoT (Chain of Thought) prompts in the SPORTU-text task, are shown in table 8 - table 11. The template prompt for the SPORTU-video multiple-choice task is shown in table 12, and for the SPORTU-video open-ended task, it is shown in Table 13.
@@ B.5 Prompt Templates
Table 8: Zero-Shot Standard Prompt for SPORTU-text
@@
content: “You are a helpful assistant who is good at answering multiple-choice sports questions.” }
@@
content: “**Question:** {{question}}. You are a sports expert who answers the above question. What are the correct answers to this question?” }
Table 9: Zero-Shot CoT for SPORTU-text
{role: “system’”,
content: “You are a helpful assistant who is good at answering multiple-choice sports questions.”}
@@
content: “**Question:** {{question}}. You are a sport expert who answers the above question. What are the correct answers for this question? Think step by step, and please explain why correct answers are correct while the rest options are incorrect. You should explain each option in detail. Add one more sentence after your detailed explanation at the end to show the correct answers’ option letters only with A, B, C, D. The correct answers: ” }
Table 10: Five-Shot Standard Prompt for SPORTU-text
@@
content: “You are a helpful assistant who is good at answering multiple-choice sports questions. Some examples are given below. {{five-shot examples}} ”}
@@
content: “**Question:** {{question}}. You are a sports expert who answers the above question. What are the correct answers to this question?”}
Table 11: Five-Shot CoT for SPORTU-text
@@
content: “You are a helpful assistant who is good at answering multiple-choice sports questions. Some examples are given below. {{five-shot examples with human-annotated explanations }}” }
{role: ‘user’,
content: “**Question:** {{question}}. You are a sports expert who answers the above question. What are the correct answers to this question?” }
Table 12: Zero-Shot Standard Prompt for SPORTU-video Multiple Choice Questions
@@
content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of the video.”}
@@
content: “**Question:** {{question}}. Answer this sports-related question based on the video{{frames}}, Respond with the letter of the correct option (A, B, C, or D), formatted as ‘The correct answer is: ’, and explain why you chose that option.”}
Table 13: Zero-Shot Standard Prompt for SPORTU-video Open-Ended Questions
@@
content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of the video.”}
@@
content: “**Question:** {{question}}. Answer this sports-related question based on the video{{frames}}. Please provide a detailed explanation focusing on the sports aspects of the video.” }
Appendix C Additional Error Analysis
@@ C.1 Error Analysis for SPORT-video
We observe that open-ended questions have the highest frequency of question understanding errors. For instance, when asked ’Why is it a foul in the video?’, the model might respond that there is no foul, which is a question understanding error because the phrasing of the question already implies that a foul occurred. This issue is much less frequent in multiple-choice questions (as shown on the left side of the figure 5). When presented with four options, the model tends to select one of the provided answers, rather than completely misinterpreting the premise of the question.
C.2 Examples of Each Error Type
Figure 6: Examples of Each Error Type
Figure 7: Examples of Each Error Type
Figure 8: Examples of Each Error Type
Figure 9: Examples of Each Error Type
Figure 10: Examples of Each Error Type
Appendix D Questions for Each Sport
@@
Appendix E Sample Questions
E.1 SPORTU-text Examples
This section shows examples of our SPORTU-text. (Table 21 - Table 26)
Table 21: Sample Question 1
@@ E.1 SPORTU-text Examples
Table 22: Sample Question 2
Question: How is a penalty awarded in football?
A) When a player intentionally handles the ball.
B) When a player scores a goal.
C) When a player passes the ball to another player.
D) When a player runs with the ball towards the goal.
Answer: A
Explanation: Handling the ball in the penalty area may result in a penalty kick, therefore option A is correct. Scoring a goal does not result in a penalty kick, therefore option B is incorrect. Passing the ball does not result in a penalty kick, therefore option C is incorrect. Dribbling towards the goal does not result in a penalty kick, therefore option D is incorrect. Hence, the correct answer is option A.
Table 23: Sample Question 3
Question: Why might a team be penalized with a loss of 5 yards in American football?
A) For a false start offense.
B) For scoring a touchdown.
C) For completing a pass.
D) For calling a timeout.
Answer: A
Explanation: A team is penalized with a loss of 5 yards in American football for a false start offense, therefore option A is correct. Option B is incorrect because it isn’t a stated rule in American football. Option C is incorrect because it isn’t a stated rule in American football. Option D is incorrect because it isn’t a stated rule in American football. Hence, the correct answer is option A.
Table 24: Sample Question 4
Question: Which of the following scores is considered a ”deuce” in a tennis match?
A) 30-30
B) 40-40
C) 15-15
D) 0-0
Answer: B
Explanation: “Deuce” refers to a score of 40-40 in tennis. 30-30 is not a deuce, therefore option A is incorrect. 40-40 is a deuce, therefore option B is correct. 15-15 is not a deuce, therefore option C is incorrect. 0-0 is not a deuce, therefore option D is incorrect. Hence, the correct answer to this question is B.
Table 25: Sample Question 5
Question: How many sets does a player need to win for a Grand Slam men’s singles tennis match to be declared the winner?
A) 1 set
B) 2 sets
C) 3 sets
D) 4 sets
@@
Explanation: Grand Slam men’s singles matches are best of five sets, meaning the first player to win three sets wins the match. Winning one set is not enough to win the match, therefore option A is incorrect. Winning two sets is also not enough to win the match, therefore option B is incorrect. The first player to win three sets wins the match, therefore option C is correct. Winning four sets is not possible since winning three sets is the maximum in a best-of-five format. Therefore, option D is incorrect. Hence, the correct answer is C.
Table 26: Sample Question 6
Question: How can the player holding the ball be considered out of bounds?
A) If the player’s foot is touching the boundary line
B) If the player is holding the ball while standing outside the boundary line
C) If the player jumps from outside the boundary line and catches the ball while in the air
D) All of the above
Answer: D
Explanation: A player is out of bounds if their foot touches the boundary line, but this is not the only factor determining if a player is out of bounds. Therefore, option A is incorrect. Standing outside the boundary line also is considered out of bounds, but again is not the only determining factor. Therefore, option B is incorrect. Catching the ball in the air while jumping from outside the boundary line also counts as out of bounds, but this is not the only determining factor. A player holding the ball is considered out of bounds if their foot touches the boundary line, if they are standing outside the boundary line while holding the ball, or if they jump from outside the boundary line and catch the ball while in the air, so all of the above answers include the factors that determine if a player is considered out of bounds. Therefore, option D is correct. Hence, the correct answer is D.
E.2 SROUTU-video Examples
This section shows examples of our SPORTU-video. (Figure 11 - Figure16)
Figure 11: Examples of SPORTU-video
Figure 12: Examples of SPORTU-video
Figure 13: Examples of SPORTU-video
Figure 14: Examples of SPORTU-video
Figure 15: Examples of SPORTU-video
Figure 16: Examples of SPORTU-video

### v3

@@ Appendix G Result of Three Prompt Settings on SPORTU-video across different levels of difficulty
           Model            Difficulty            Performance            X-YR            X-RY            X-Y            Claude-3.0-Haiku            Easy            68.41            66.58            66.62            Medium            46.43            46.11            46.53            Hard            20.12            18.94            22.42            Overall            48.07            47.19            47.95            Claude-3.5-Sonnet            Easy            88.65            63.83            89.15            Medium            65.10            55.52            65.88            Hard            52.57            39.32            53.06            Overall            69.52            55.08            70.18            Gemini 1.5 Pro            Easy            85.85            85.20            87.52            Medium            58.25            58.11            61.22            Hard            43.53            42.75            39.98            Overall            65.13            63.04            64.93            Gemini 1.5 Flash            Easy            85.99            59.07            85.19            Medium            53.38            50.85            58.56            Hard            38.73            12.89            38.26            Overall            59.97            46.68            62.52            GPT-4omini            Easy            66.09            59.02            66.68            Medium            55.69            42.25            58.12            Hard            47.54            13.67            44.49            Overall            57.24            42.06            58.19            GPT-4o            Easy            84.89            79.92            84.30            Medium            62.31            65.98            64.83            Hard            57.84            40.51            56.20            Overall            68.00            65.56            68.79            Qwen2-VL-72B            Easy            94.86            84.16            95.11            Medium            66.27            61.69            66.97            Hard            36.53            30.56            44.12            Overall            69.18            62.65            70.94            ChatUniVi            Easy            59.22            49.04            55.99            Medium            36.95            28.55            35.63            Hard            32.21            18.71            39.07            Overall            42.35            32.58            41.89            LLaVA-NeXT            Easy            94.24            91.43            92.44            Medium            67.02            56.00            59.39            Hard            33.44            34.21            30.78            Overall            68.89            62.16            63.72            mPLUG-Owl3            Easy            87.28            88.51            87.11            Medium            55.58            58.75            57.37            Hard            25.40            24.88            28.89            Overall            59.26            61.27            60.80            ST-LLM            Easy            59.26            68.54            63.09            Medium            36.08            33.23            41.14            Hard            30.51            22.10            36.08            Overall            41.59            40.09            46.39            Tarsier            Easy            89.71            84.01            88.13            Medium            58.43            50.94            58.35            Hard            24.25            24.80            25.17            Overall            61.32            55.70            60.99            Video-ChatGPT            Easy            62.36            63.91            37.04            Medium            39.40            38.64            36.01            Hard            32.55            19.12            22.82            Overall            44.63            42.36            34.05            VideoChat2            Easy            88.30            88.62            89.37            Medium            58.73            60.45            58.65            Hard            26.94            28.13            25.29            Overall            61.55            62.79            61.53
Appendix H Examples of Errors Across Prompt Strategies
In this section, we show examples where the question was answered correctly in the X $\rightarrow Y$ prompt setting but failed in either the X $\rightarrow RY$, X $\rightarrow YR$, or both settings.
Figure 6: Basketball Easy level Question
Figure 7: The same question is answered correctly in the $X\rightarrow Y$ setting but incorrectly in both the $X\rightarrow RY$ and $X\rightarrow YR$ settings.
Figure 8: The same question is answered correctly in both the $X\rightarrow Y$ setting and $X\rightarrow YR$ settings but incorrectly in the $X\rightarrow RY$ setting.
Appendix I Correlation Matrix of Evaluation Metrics
@@
We applied Pearson correlation to assess the relationship between the automatic evaluation metrics and human-annotated scores for the SPORTU-Video part. As shown in Figure 9, G-eval has the highest correlation with the human scores among all the automatic metrics. However, the correlation is still weak, indicating the need for the development of a new evaluation metric in the future to improve the assessment process.
Figure 9: Examples of SPORTU-video
Appendix J Criteira of G-Eval
Figure 10: Criteria and prompt used in G-Eval score evaluation. The same prompt template is applied to all four evaluator models.
Appendix K SPORTU-text Prompt Templates
K.1 zero shot standard prompt
@@
Table 12: Prompt Template for Zero-shot Standard Prompt on SPORTU-text
@@
 content: “You are a sport experts answering sports-related questions. Please indicate the correct answer(s) clearly with letter(s).”
@@
 content: “Question: {{question}} {{options}}. Only output the correct option letters.”
K.2 zero shot CoT prompt
Table 13: Prompt Template for Zero-shot CoT Prompt on SPORTU-text
 role: “system”,
 content: “You are a sport experts answering sports-related questions. Please indicate the correct answer(s) clearly with letter(s).”
@@
 content: “ Question: {{question}} {{options}}. Let’s think step by step.”
K.3 five shot standard prompt
Table 14: Prompt Template for Five-shot Standard Prompt on SPORTU-text
@@
 content: “You are a sport experts answering sports-related questions. Please indicate the correct answer(s) clearly with letter(s).”
{{five-shot examples}}, {
@@
 content: “Question: {{question}} {{options}}. Only output the correct option letters.”
K.4 Five shot CoT prompt
Table 15: Prompt Template for Five-shot CoT Prompt on SPORTU-text
@@
 content: “You are a sport experts answering sports-related questions. Please indicate the correct answer(s) clearly with letter(s).”
{{five-shot examples with explanations}}, {
 role:“user”,
 content: “Question: {{question}} Options: {{options}}.”
Appendix L SPORTU-video Prompt Templates
L.1 X - YR Prompt Template
Table 16: Prompt Template for SPORTU-video $X\rightarrow YR$ setting
@@
 content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of these frames.”
@@
 content: [
  {{video_frames}},
   type: “text”,
   text: “Based on the video frames provided, answer this sports-related question: {{question}} Options: {{options}}. Respond with the letter of the correct option, formatted as ‘The correct answer is: ’, and explain why you chose that option.”
L.2 X - RY Prompt Template
Table 17: Prompt Template for SPORTU-video $X\rightarrow RY$ setting
@@
 content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of these frames.”
 role: “user”
 content: [
  {{video_frames}},
   type: “text”,
   text: “Based on the video frames provided, answer this sports-related question: {{question}} Options: {{options}}. Let’s answer this question step by step. add a sentence formatted as ’The correct answer is: ’ at the end of your thinking process. ”
L.3 X - Y Prompt Template
Table 18: Prompt Template for SPORTU-video $X\rightarrow Y$ setting
 role: “system”,
 content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of these frames.”
@@
 content: [
  {{video_frames}},
   type: “text”,
   text: “Based on the video frames provided, answer this sports-related question: {{question}} Options: {{options}}. Answer with only the letter of the correct option.”
L.4 Open-ended Template
Table 19: Prompt Template for SPORTU-video Open-ended Questions
 role: “system”,
 content: “You are a sports expert analyzing a series of video frames that form a continuous short video clip. Your task is to answer questions based on the content of these frames.”
 role: “user”,
 content: [
  {{video_frames}},
   type: “text”,
   text: “Based on the video frames provided, answer this sports-related question: {{question}}. Please provide a detailed explanation, focusing on the sports aspects of the video.”
Appendix M Additional Error Analysis
@@
We observe that open-ended questions have the highest frequency of question understanding errors. For instance, when asked ‘Why is it a foul in the video?’, the model might respond that there is no foul, which is a question understanding error because the phrasing of the question already implies that a foul occurred. This issue is much less frequent in multiple-choice questions (as shown on the left side of the figure 11). When presented with four options, the model tends to select one of the provided answers, rather than completely misinterpreting the premise of the question. Examples of each Error are in Appendix Q.
Appendix N Question Template for Each Sport
@@
Appendix O SPORTU-text Examples
O.1 Rule-related Question
Table 27: Rule Question 1
Question: How does ball possession work in basketball after the successful final free throw?
A) The team that made the free throw retains possession.
B) The team that missed the free throw gains possession.
C) The opposing team gains possession.
D) The team with the most points gains possession.
Answer: C
Explanation: According to FIBA’s rules on what should happen after a successful field goal or the last successful free throw, following the last successful free throw, any player from the non-scoring team should inbound the ball from any position behind their own endline; therefore, option A is incorrect. The team that missed the free throw is the same team as the one executing the free throw during the game, but the question specifies that the team’s last free throw was successful; therefore, option B is incorrect. Since the rule states that any player from the non-scoring team should inbound the ball from behind their own endline, and since the team that executed the last free throw scored, the non-scoring team, i.e., the opposing team, should regain possession of the ball and inbound it; therefore, option C is correct. The scoring situation of both teams does not affect the distribution of possession after a successful free throw, so neither the number of points scored nor the point differential is a determining factor for possession distribution; therefore, option D is incorrect. Hence, the correct answer is option C.
Table 28: Rule Question 2
Question: Question: Which player is typically responsible for throwing the ball to the receivers in American football?
A) Linebacker.
B) Quarterback.
C) Running back.
D) Tight end.
Answer: B
Explanation: Linebacker is on the defensive side and does not often have a chance to touch the ball, therefore option A is incorrect. Quarterback usually throws the ball to running back, therefore option B is correct. Running back is the one who often receives the ball, therefore option C is incorrect. Tight end usually blocks the other player and is not primarily meant to grab the ball, therefore option D is incorrect. Hence, the correct answer is option B.
O.2 Strategy-related Question
Table 29: Strategy-related Question Sample 1
Question: What are key considerations when implementing a successful blocking scheme in volleyball defensive strategies?
A) The blocker’s positioning in relation to the attacker’s hitting arm.
B) The speed of the incoming serve.
C) The position of the setter on the opposing team.
D) The timing and coordination between the blockers.
Answer: ABCD
Explanation: The positioning in relation to the attacker’s hitting arm is crucial for a successful block and for the entire team’s block and defensive formation. The side blocker usually positions their head in line with the hitter’s arm and the ball to cover the straight line, or uses their right hand on the hitter’s arm and the ball to cover the cross-court angle. The back-row players will adjust their position based on the main angles that the blockers are covering. Therefore, option A is correct. Increasing the speed of the serve makes it harder for the opponent to receive, potentially putting them out of system. This reduces the opponent’s attacking options compared to when they are in system, making blocking easier. Therefore, option B is correct. The setter’s position directly determines the team’s strategic options and makes it easier for blockers to read and predict the play. Therefore, option C is correct. Timing and coordination between the blockers are also essential, as blockers aim to jump together without leaving gaps if they decide to execute a multi-blocker jump. Therefore, option D is correct. Hence, the correct answer to this question is A, B, C, and D.
Table 30: Strategy-related Question Sample 2
Question: Question: Which of the following are common attack patterns in volleyball offensive strategies?
A) Quick set
B) Slide attack
C) 5-1 formation
D) Triple quick
Answer: AB
Explanation: A quick set is an attacking pattern for the middle blocker. The middle blocker jumps close to the setter and jumps before the setter sets the ball. The setter then sets the ball low and quickly to the middle blocker’s optimal attacking position. Therefore, option A is correct. A slide attack is another attacking pattern for the middle blocker, where the middle blocker approaches the right-side position, usually 2-3 meters from the setter, and swings with a one-foot jump. The setter will set a low arc ball backward to the right side. Therefore, option B is correct. The 5-1 formation is a team strategy concerning how many setters are on the court and is not related to attack patterns. Therefore, option C is incorrect. ‘̀Triple quickïs not a common attack pattern or strategy in standard volleyball. Therefore, option D is incorrect. Hence, the correct answers are A and B.
O.3 Scenario-related Question
Table 31: Scenario-Related Question 1
@@
Table 32: Scenario-Related Question 2
Question: In a competitive basketball game, Player A accidentally knocks the ball which then bounces off Player B’s hand, rolls on the boundary line, touches Player C’s foot while he is standing on the line, and finally goes out of the court. Which player is considered to have caused the ball to go out of bounds?
A) A) Player A because he accidentally knocked the ball first.
B) Player B because the ball touched his hand before rolling on the boundary line.
C) Player C because the ball touched his foot while he was standing on the boundary line.
D) None of the players caused the ball to go out of bounds because the ball rolled on its own.
@@
Explanation: According to FIBA’s rules regarding out-of-bounds, only the player who last touched or was touched by the ball before it went out of bounds is considered responsible for the out-of-bounds situation. Since Player A touched the ball before it touched Player B, Player A was not the last player to touch the ball; therefore, option A is incorrect. Similarly, after Player B touched the ball, it then touched Player C, so Player B was not the last player to touch the ball; therefore, option B is incorrect. Because Player C was the last player to touch the ball before it was deemed out of bounds, and the ball touched Player C’s foot, causing it to go out of bounds, Player C is considered responsible for the out-of-bounds situation; therefore, option C is correct. Since the ball touched at least one player before going out of bounds, it cannot be considered that no player caused the ball to go out of bounds; therefore, option D is incorrect. Hence, the correct answer is option C.
Appendix P SROUTU-video Examples
P.1 Basketball
Figure 12: Basketball Easy level Question
Figure 13: Basketball Medium level Question
Figure 14: Basketball hard level Question
P.2 Volleyball
Figure 15: Volleyball easy level Question
Figure 16: Volleyball medium level Question
Figure 17: Volleyball hard level Question
P.3 Soccer
Figure 18: Soccer easy level Question
Figure 19: Soccer medium level Question
Figure 20: Soccer hard level Question
P.4 Badminton
Figure 21: Badminton easy level Question
Figure 22: Badminton medium level Question
Figure 23: Badminton hard level Question
P.5 American Football
Figure 24: American Football easy level Question
Figure 25: American Football medium level Question
Figure 26: American Football hard level Question
P.6 Ice Hockey
Figure 27: Ice Hockey easy level Question
Figure 28: Ice Hockey medium level Question
Figure 29: Ice Hockey hard level Question
P.7 Baseball
Figure 30: Baseball easy level Question
Figure 31: Baseball medium level Question
Figure 32: Baseball hard level Question
Appendix Q Examples of Each Error Type
Q.1 Question Understanding Error
Figure 33: Examples of Question Understanding Type and Error Reason Explanations
Figure 34: Examples of Question Understanding Error and Error Reason Explanation
Figure 35: Examples of Question Understanding Error and Error Reason Explanation
Q.2 Visual Perception Error
Figure 36: Examples of Visual Perception Error and Error Reason Explanation
Figure 37: Examples of Visual Perception Error and Error Reason Explanation
Figure 38: Examples of Visual Perception Error and Error Reason Explanation
Q.3 Hallucination Error
Figure 39: Examples of Hallucinations Error and Error Reason Explanation
Figure 40: Examples of Hallucination Error and Error Reason Explanation
Figure 41: Examples of Hallucination Error and Error Reason Explanation
Q.4 Reasoning Error
Figure 42: Examples of Reasoning Error and Error Reason Explanation
Figure 43: Examples of Reasoning Error and Error Reason Explanation
Figure 44: Examples of Reasoning Error and Error Reason Explanation
Q.5 Lack of Domain Knowledge
Figure 45: Examples of Lack of Domain Knowledge and Error Reason Explanation
Figure 46: Examples of Lack of Domain Knowledge and Error Reason Explanation
Figure 47: Examples of Lack of Domain Knowledge and Error Reason Explanation